OpenStudy (thehollywoodstud):

Need Help @janu16 will Medal you

OpenStudy (thehollywoodstud):

@Janu16

jhonyy9 (jhonyy9):

write the proportionalitie of corresponding sides

OpenStudy (janu16):

ok

jhonyy9 (jhonyy9):

to side with length 8 corespond the side with length 2 so to side x-3 coresp. the side with length 2,5 cross multiplie and calcul the x = ?

OpenStudy (thehollywoodstud):

ok i solved it thx

jhonyy9 (jhonyy9):

ok. good luck

OpenStudy (janu16):

Smaller one- 2.5*4= 10 2*4= 8 4*4= 16 x-3=10 so x =13